What is always the different between two isotopes on an element? (Choose all correct answers)

number of electrons

number of protons

number of neutrons
Answer #1

Ans :- Option (c) " Number of neutrons " is always the different between two isotopes on an element.

Explanation :-

Because, Isotopes have not only the same atomic number but also have same number of protons and electrons. The number of neutrons and mass number are not same in isotopes.

For example :- Three isotopes of Hydrogen have same atomic number, same number of protons and electrons but do not have same number of neutrons.

1). Protium (1H1) :- Atomic number = Number of protons =Number of electrons = 1 but Number of neutrons = 1-1 = 0

2). Deuterium (1H2) :- Atomic number = Number of protons =Number of electrons = 1 but Number of neutrons = 2-1 = 1

3). Tritium (1H3) :- Atomic number = Number of protons =Number of electrons = 1 but Number of neutrons = 3-1 = 2

Hence, Number of neutrons is the correct answer.